OpenStudy (anonymous):

The Coefficient of kinetic friction between the tires of your car and the roadway is "μ". (a) If your initial speed is "v" and you lock your tires during breaking, how far do you skid? (b) What is the stopping distance for a truck with twice the mass of you car, assuming same initial speed and coefficient of kinetic friction. I'm stuck on (b), my mass cancels out every time, but that doesn't seem right

OpenStudy (amistre64):

im thinking its 4 times as much, but i cant recall the coeff of friction stuff too well. can you show your process?

OpenStudy (amistre64):

Force = coeff * normal .... maybe

OpenStudy (amistre64):

does this look familiar? cant tell if im remembering it correctl \[d_{stop}=\frac{\frac{1}{2}mv^2}{\mu_k~\vec n}\]

OpenStudy (anonymous):

Force = coefficient* normal reaction= (u_k)*mg Acceleration (or retardation in this case), a = force/ mass= (u_k)*g stopping distance can be found from : v^2=u^2-2aS (where v=0) That is, S= u^2/2*a In either case the distance is is independent of the truck or the car as the acceleration is independent of the mass.